OpenStudy (anonymous):

x-3y=9.. I've tried isolating y but it's gotten to the point i confused myself and have gotten two different answers, my last being y=x-9 over -3

Nnesha (nnesha):

first step *)subtract x both side

Nnesha (nnesha):

 what you should get when you subtract x both side what sign suppose to be at right side ?

OpenStudy (anonymous):

\[x-3y=9\]\[x-9=3y\]\[3y=x-9\] Now divide both x and -9 by 3. You should get\[y=\frac{ 1 }{ 3 }x-3\]as your answer.

OpenStudy (anonymous):

@nnesha, i'm still confused because that's what i did and im still getting the same answer.. this is what ive been doing